# An Efficient Numerical Scheme for a Time-Fractional Burgers Equation with Caputo-Prabhakar Derivative

### v1

@@ 3.1. Newton’s iterative method
$\psi(u^{(s+1)})=\psi(u^{(s)})+(u^{(s+1)}-u^{(s)})\psi^{\prime}(u^{(s)}).$ (3.3)
Whence Eq. (3.1) at $(s+1)^{th}$ level becomes
@@
Now, after applying the formula (3.3) we get the following equation
@@
$\delta^{2}_{x}(u^{(s+1)})^{k}_{i}$ $=\frac{1}{h^{2}}[(u^{(s+1)})^{k}_{i+1}-2(u^{(s+1)})^{k}_{i}+(u^{(s+1)})^{k}_{i-1}],$ (3.5)
$\psi(u^{(s)})^{k}_{i}$ $=\frac{1}{6h}[(u^{(s)})^{k}_{i+1}+(u^{(s)})^{k}_{i}+(u^{(s)})^{k}_{i-1}][(u^{(s)})^{k}_{i+1}-(u^{(s)})^{k}_{i-1}],$ (3.6)
$\psi^{\prime}(u^{(s)})^{k}_{i}$ $=\frac{1}{2h}[(u^{(s)})^{k}_{i+1}-(u^{(s)})^{k}_{i-1}].$ (3.7)
So after putting the values (3.5)-(3.7) and doing some simplification, the vector representation of Eq. (3.4) becomes
@@
$=\frac{\mu}{2h}\text{diag}(E_{2}.(U^{(s)})^{k})(U^{(s)})^{k}$ $-\frac{\mu}{6h}\text{diag}(E_{3}.(U^{(s)})^{k})E_{2}(U^{(s)})^{k}+H^{k},$ (3.8)
@@
$E_{3}$ $=\begin{bmatrix}1&1&0&&\\ 1&1&1&&\\ &\ddots&\ddots&\ddots&\\ &&1&1&1\\ &&0&1&1\end{bmatrix}_{M-1\times M-1};~{}~{}~{}~{}U^{k}=\begin{bmatrix}u_{1}^{k}\\ u_{2}^{k}\\ \cdots\\ u_{M-1}^{k}\end{bmatrix};~{}~{}~{}~{}F^{k}=\begin{bmatrix}f_{1}^{k}\\ f_{2}^{k}\\ \cdots\\ f_{M-1}^{k}\end{bmatrix}.$
@@
$-\left.\frac{\mu}{6h}\text{diag}(E_{3}.(U^{(s)})^{k})E_{2}(U^{(s)})^{k}+H^{k}\right].$ (3.9)
$\Xi_{j}=\max_{\begin{subarray}{c}{0<i<N}\\ {0<k<M}\end{subarray}}$ $|\mathcal{U}(x_{i},t_{k})-u^{k}_{i}|,~{}~{}~{}~{}(\text{at}~{}j^{\text{th}}~{}\text{iteration,})$, ) (3.10)
$\Theta_{j+1}=$ $\ \log_{2}\left(\frac{\Xi_{j}}{\Xi_{j+1}}\right),$ (3.11)
Example 1.
Consider the problem (1.2)–(1.5) with $u_{0}(x)=0$ and
The exact solution of (1.2)–(1.5) with the above conditions is $\ \mathcal{U}(x,t)=t^{2}\sin(\pi x)$. To evaluate the numerical solution of this example, we choose Maximum Step = 500 (see Appendix A). We have two tables to show the maximum error ($\Xi$) and convergence order($\Theta$) of the scheme. Table 1 is in space direction where iterative accuracy ($ItAcc$) is E-05, and Table 2 is in temporal direction with $ItAcc$ = E-08. In both tables, we found that the order of convergence is nearly two as the convergence order of the presented numerical scheme is $(\tau^{2-\alpha}+h^{2})$.
Table 1. The error $\Xi$, order $\Theta$, computing time, and number of iterations for Example 1 with $N=2^{14}$, eps=1e-5, $MaxStep$ = 500.
@@
Table 2. The error $\Xi$, order $\Theta$, computing time, and number of iterations for Example 1 with $M=2^{12},$ $eps=1e-8$, $MaxStep=500$.
@@
Consider Eq.(1.2)-(1.5) with $u_{0}(x)=0$. Let $L=1=T=1,~{}~{}f(x,t)=120x(x-1)t^{5-\alpha}E_{\rho,6-\alpha}^{-\gamma}(\omega t^{\rho})+x(x-1)(2x-1)t^{10}-2t^{5}$.
The exact solution of this example is $\ \mathcal{U}(x,t)=t^{5}x(x-1)$. Here, we choose iterative accuracy $ItAcc$ = E-08 for time direction (Table 3) and $ItAcc$ = E-06 for space direction (Table 4), and Maximum Step = 500 (see Appendix A). Again, the obtained order of convergence of the scheme is approximately two.
Table 3. The error $\Xi$, order $\Theta$, computing time, and number of iterations for Example 2 with $M=2^{11}$, $MaxStep=500$, $eps=1e-8$.
@@
Table 4. The error $\Xi$, order $\Theta$, computing time, and number of iterations for Example 2 with $N=2^{15}$, $eps=1e-6$, $MaxStep=500$.
@@
4. Conclusions
In conclusion, we have effectively solved the fractional Burgers equation using an implicit difference scheme that incorporates the Caputo-Prabhakar derivative. Our analysis establishes the stability and convergence of the proposed numerical scheme, demonstrating a convergence order of $(\tau^{2-\alpha}+h^{2})$. The method described in this paper achieves a better order of convergence in time direction, as compared to the paper [35]. Moreover, the numerical experiments conducted support our theoretical findings, illustrating the robustness and effectiveness of our approach.
Acknowledgments
Authors thanks W. Qiu for the help when the work was at initial level. DS is supported by UGC under SRF scheme. SY is supported by MUR (PRIN2022 research grant n. 202292JW3F). This work is partially supported by the ERCIM ‘Alain Bensoussan’ Fellowship programme.

### v2

@@ 3.1 Newton’s iterative method
$\psi(u^{(s+1)})=\psi(u^{(s)})+(u^{(s+1)}-u^{(s)})\psi^{\prime}(u^{(s)}).$ (24)
Whence Eq. (22) at $(s+1)^{th}$ level becomes
@@
Now, after applying the formula (24) we get the following equation
@@
$\delta^{2}_{x}(u^{(s+1)})^{k}_{i}$ $=\frac{1}{h^{2}}[(u^{(s+1)})^{k}_{i+1}-2(u^{(s+1)})^{k}_{i}+(u^{(s+1)})^{k}_{i-1}],$ (26)
$\psi(u^{(s)})^{k}_{i}$ $=\frac{1}{6h}[(u^{(s)})^{k}_{i+1}+(u^{(s)})^{k}_{i}+(u^{(s)})^{k}_{i-1}][(u^{(s)})^{k}_{i+1}-(u^{(s)})^{k}_{i-1}],$ (27)
$\psi^{\prime}(u^{(s)})^{k}_{i}$ $=\frac{1}{2h}[(u^{(s)})^{k}_{i+1}-(u^{(s)})^{k}_{i-1}].$ (28)
So after putting the values (26)-(28) and doing some simplification, the vector representation of Eq. (25) becomes
@@
$=\frac{\mu}{2h}\text{diag}(E_{2}.(U^{(s)})^{k})(U^{(s)})^{k}$ $-\frac{\mu}{6h}\text{diag}(E_{3}.(U^{(s)})^{k})E_{2}(U^{(s)})^{k}+H^{k},$ (29)
@@
$E_{3}$ $=\begin{bmatrix}1&1&0&&\\ 1&1&1&&\\ &\ddots&\ddots&\ddots&\\ &&1&1&1\\ &&0&1&1\end{bmatrix}_{M-1\times M-1};~~~~U^{k}=\begin{bmatrix}u_{1}^{k}\\ u_{2}^{k}\\ \cdots\\ u_{M-1}^{k}\end{bmatrix},~~~~F^{k}=\begin{bmatrix}f_{1}^{k}\\ f_{2}^{k}\\ \cdots\\ f_{M-1}^{k}\end{bmatrix}$.
The operator $\text{diag}(.)$ transforms the vector into a diagonal matrix. We substitute $B:=\left[a_{1}I-\frac{\mu}{h^{2}}E_{1}+\frac{\mu}{2h}E_{2}.(U^{(s)})^{k}\right]$ in (3.1). The matrix $B$ is invertible, hence we obtain the final scheme as
$-\left.\frac{\mu}{6h}\text{diag}(E_{3}.(U^{(s)})^{k})E_{2}(U^{(s)})^{k}+H^{k}\right].$ (30)
$\Xi_{j}:=\max_{\begin{subarray}{c}{0<i<N}\\ {0<k<M}\end{subarray}}$ $|\mathcal{U}(x_{i},t_{k})-u^{k}_{i}|,~~~~(\text{at}~j^{\text{th}}~\text{iteration,})$ (31)
$\Theta_{j+1}:=$ $\ \log_{2}\left(\frac{\Xi_{j}}{\Xi_{j+1}}\right).$ (32)
Example 1.
Consider the problem (2)–(5) with $u_{0}(x)=0$ and
@@
Table 1: The error $\Xi$, order $\Theta$, computing time, and number of iterations for Example 1 with $N=2^{14}$, ItAcc = 1E-05, MaxStep = 500.
@@
Table 2: The error $\Xi$, order $\Theta$, computing time, and number of iterations for Example 1 with $M=2^{12},$ ItAcc=1E08ItAcc=1E-08, $MaxStep=500$.
@@
Consider Eq.(2)-(5) with $u_{0}(x)=0$. Let $L=1=T=1,~~f(x,t)=120x(x-1)t^{5-\alpha}E_{\rho,6-\alpha}^{-\gamma}(\omega t^{\rho})+x(x-1)(2x-1)t^{10}-2t^{5}$.
[…]
Table 3: The error $\Xi$, order $\Theta$, computing time, and number of iterations for Example 2 with $M=2^{11}$, $MaxStep=500$, ItAcc=1E08ItAcc=1E-08.
@@
Table 4: The error $\Xi$, order $\Theta$, computing time, and number of iterations for Example 2 with $N=2^{15}$, ItAcc=1e6ItAcc=1e-6, $MaxStep=500$.
@@
4 Conclusions
In conclusion, we have effectively solved the fractional Burgers equation using an implicit difference scheme that incorporates the Caputo-Prabhakar derivative. Our analysis establishes the stability and convergence of the proposed numerical scheme, demonstrating a convergence order of $(\tau^{2-\alpha}+h^{2})$. The method described in this paper achieves a better order of convergence in time direction, as compared to the paper [37]. Moreover, the numerical experiments conducted support our theoretical findings, illustrating the robustness and effectiveness of our approach. While this work establishes a solid foundation for the numerical analysis of the fractional Burgers equation, we have not considered the initial singularity of the solution, which is commonly observed in many fractional PDEs [20, 35]. Addressing this aspect will be the focus of our future study. However, the approach discussed here still provides a strong basis for future advancements and remains applicable in cases where the initial singularity has minimal impact or can be addressed through alternative methods in subsequent analysis.
[…]
Acknowledgements Authors thanks W. Qiu for the help when the work was at initial level.
Declarations
  • Funding: DS is supported by UGC under SRF scheme. SY is supported by MUR (PRIN2022 research grant n. 202292JW3F). This work is partially supported by the ERCIM ‘Alain Bensoussan’ Fellowship programme.
  • Conflict of interest/Competing interests: The authors have no competing interests to declare.
  • Ethics approval and consent to participate: Not applicable.
  • Consent for publication: Not applicable.
  • Data availability: Not applicable.
  • Materials availability: Not applicable.
  • Code availability: Not applicable.
  • Author contribution: All the authors contributed equally to this work.